# Model Human Learners: Computational Models to Guide Instructional Design

### v3

@@ Introduction
In their seminal paper, ? (?) propose the concept of a Model Human Processor, a unified information-processing model that codifies psychological theories to predict human performance in user interfaces. They argue that interface designers could use this kind of model to make more informed design decisions that improve usability. In this paper, I propose an analogous concept: a Model Human Learner, a unified computational model that codifies cognitive and learning theories to predict human learning in instructional systems. I argue that instructional designers could use this kind of model to make more informed instructional design decisions that improve pedagogical effectiveness.
@@
An analysis by ? (?) suggests that instructional designers face a crisis of choice—design spaces are combinatorial, containing trillions of possible configurations. How can instructional designer determine the best choices? Currently, they rely on A/B experiments to navigate this vast space, but human studies are costly and time consuming. Moreover, each A/B experiment provides only a single bit of information, meaning instructional designers are essentially playing a losing game of 20 questions with nature (?, ?).
@@
Computational models of learning, such as those proposed by ? (?) and ? (?), offer a potential solution. Unlike abstract, mathematical models of learning, such as the Additive Factors Model (?, ?) or Bayesian Knowledge Tracing (?, ?), which fit functions to performance data, computational models of learning are mechanistic. Similar to cognitive architectures (?, ?), these models use artificial intelligence and machine learning to simulate how knowledge is updated in response to practice and how performance evolves over time. I argue that these models can realize the Model Human Learner concept, letting instructional designers simulate A/B experiments to test alternative interventions and identify the most promising ones before conducting costly human studies.
@@
Prior research has explored several applications of computational models of learning. For example, ? (?) investigated their use for discovering cognitive models, while ? (?) explored how they can promote “learning-by-teaching,” where students learn by instructing simulated students. ? (?) and ? (?) examined their applications for tutor authoring. Lastly, others have studied their utility for theory testing (?, ?, ?, ?).
@@
More recently, researchers have proposed using these models to predict how different instructional design choices will impact learning (?, ?, ?). While these studies have generated reasonable predictions about experimental effects, none have yet been compared with human outcomes. This critical next step is the focus of this paper. Specifically, I apply a computational model of learning to predict the outcomes of two human A/B experiments. By comparing these predictions to previously collected human data, I demonstrate that computational models of learning can:
@@
  • Successfully predict the main effects of two human experiments—one evaluating a problem sequencing intervention and the other testing an item design intervention;
@@
  • Generate learning curve predictions that closely align with human learning curve trends, without training on human data first; and
@@
  • Provide theoretical insights into why specific interventions work, challenging a prior hypothesis by ? (?) and suggesting a novel explanation.
@@ Conclusion
To my knowledge, this paper presents the first evidence that a computational model of learning can successfully predict the main effects observed in multiple human A/B experiments. It also demonstrates that this models can generate reasonable predictions of human learning trajectories and offer theoretical insights into the effectiveness of specific instructional interventions. These findings suggest that computational models of learning could operate as Model Human Learners, supporting instructional designers in evaluating and identifying the most pedagogically effective instructional designs. More research is needed to expand the Model Human Learner concept to additional tasks and interventions, but this work represents a preliminary proof of concept that lays the groundwork for a broader research program.
@@ Acknowledgments
This work grew directly out of my PhD thesis (?, ?). As a result, it would not have been possible without the members of my thesis dissertation committee (Ken Koedinger, Pat Langley, Vincent Aleven, and John Anderson) who provided me with invaluable feedback, guidance, and support. This work has been generously funded in part from several sources, including a Graduate Training Grant awarded to Carnegie Mellon University by the Department of Education (#R305B090023 and #R305A090519), by the Pittsburgh Science of Learning Center, which is funded by the NSF (#SBE-0836012), two National Science Foundation Awards (#DRL-0910176 and #DRL-1252440), a DARPA award (#HR00111990055), and the NSF National AI Institutes Program (#2247790 and #2112532). The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. I would also like to thank Rony Patel, Hee Seung Lee, and the PSLC DataShop team for providing access to the data used in this paper. Finally, this work would also not have been possible without my collaborator Erik Harpstead, who was integral in first articulating the idea of a computational model of human learning in our original paper on the topic (?, ?).

### v4

@@ Introduction
In their seminal paper, ? (?) propose the concept of a Model Human Processor—a unified information-processing model that codifies psychological theories to predict human performance in user interfaces. They argue that interface designers could use such a model to make more informed design decisions that improve usability. I propose an analogous concept: a Model Human Learner—a unified computational model that codifies cognitive and learning theories to predict human learning in instructional systems. I argue that instructional designers could use this kind of model to make more informed instructional design decisions that improve pedagogical effectiveness.
@@
An analysis by ? (?) suggests that instructional designers face a crisis of choice, as design spaces are combinatorial and contain trillions of configurations. How can instructional designer determine the best choices? Currently, they rely on A/B experiments to navigate this vast space, but human studies are costly and time consuming. Also, each experiment provides only a single bit of information, so instructional designers are essentially playing a losing game of 20 questions with nature (?, ?).
@@
Computational models of learning, such as those proposed by ? (?) and ? (?), offer a potential solution. Unlike abstract, mathematical models of learning, such as the Additive Factors Model (?, ?) or Bayesian Knowledge Tracing (?, ?), which fit functions to performance data, computational models of learning are mechanistic. Similar to cognitive architectures (?, ?), these models simulate how knowledge is updated in response to practice and how performance evolves over time. I argue that these models can realize the Model Human Learner concept, letting instructional designers simulate A/B experiments to test alternative interventions and identify the most promising ones before conducting costly human studies.
@@
Prior research has explored several applications of computational models of learning. ? (?) investigated their use for discovering cognitive models, while ? (?) explored how they can promote “learning-by-teaching,” where students learn by instructing simulated students. ? (?) and ? (?, ?) examined their applications for tutor authoring. Others have studied their utility for theory testing (?, ?, ?, ?).
@@
Researchers have also proposed using these models to predict the effects of different instructional design choices (?, ?, ?). While these studies generate reasonable predictions, they have not yet been validated against human data. Some prior work has examined model-human alignment (?, ?, ?, ?, ?), but this has focused primarily on theory testing, rather than instructional design evaluation. This paper addresses that gap by comparing model predictions to outcomes from two human A/B experiments. Specifically, I demonstrate that computational models of learning can:
@@
  • Accurately predict the main effects of two human experiments—one evaluating a problem sequencing intervention and the other testing an item design intervention;
@@
  • Generate learning curve predictions that closely match human trends, without training on human data first; and
@@
  • Offer theoretical insights into why specific interventions work, challenging a prior hypothesis by ? (?) and suggesting a novel explanation.
@@ Conclusion
To my knowledge, this paper presents the first evidence that a computational model of learning can successfully predict the main effects observed in multiple human A/B experiments. It also demonstrates that this model can generate reasonable predictions of human learning trajectories and offer theoretical insights into the effectiveness of specific instructional interventions. These findings suggest that computational models of learning could operate as Model Human Learners, supporting instructional designers in evaluating and identifying the most pedagogically effective instructional designs. More research is needed to expand the Model Human Learner concept to additional tasks and interventions, but this work represents a preliminary proof of concept that lays the groundwork for a broader research program.
@@ Acknowledgments
This work grew out of my PhD dissertation (?, ?). As a result, it would not have been possible without the members of my thesis committee (Ken Koedinger, Pat Langley, Vincent Aleven, and John Anderson) who provided me with invaluable feedback, guidance, and support. This work has been generously funded in part from several sources, including a Graduate Training Grant awarded to Carnegie Mellon University by the Department of Education (#R305B090023 and #R305A090519), by the Pittsburgh Science of Learning Center, which is funded by the NSF (#SBE-0836012), two National Science Foundation Awards (#DRL-0910176 and #DRL-1252440), a DARPA award (#HR00111990055), and the NSF National AI Institutes Program (#2247790 and #2112532). The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. I would also like to thank Rony Patel, Hee Seung Lee, and the PSLC DataShop team for providing access to the data used in this paper. Finally, this work would also not have been possible without my collaborator Erik Harpstead, who collaborated with me to first articulate the idea of a computational model of human learning in our original paper on the topic (?, ?).